### v3

# Link Prediction with Physics-Inspired

Abstract
The message-passing mechanism underlying Graph Neural Networks (GNNs) is not naturally suited for heterophilic datasets, where adjacent nodes often have different labels. Most solutions to this problem remain confined to the task of node classification. In this article, we focus on the valuable task of link prediction under heterophily, an interesting problem for recommendation systems, social network analysis, and other applications. GNNs like GRAFF have improved node classification under heterophily by incorporating physics biases in the architecture. Similarly, we propose GRAFF-LP, an extension of GRAFF for link prediction. We show that GRAFF-LP effectively discriminates existing from non-existing edges by learning implicitly to separate the edge gradients. Based on this information, we propose a new readout function inspired by physics. Remarkably, this new function not only enhances the performance of GRAFF-LP but also improves that of other baseline models, leading us to reconsider how every link prediction experiment has been conducted so far. Finally, we provide evidence that even simple GNNs did not experience greater difficulty in predicting heterophilic links compared to homophilic ones. This leads us to believe in the necessity for heterophily measures specifically tailored for link prediction, distinct from those used in node classification. The code and appendix are available at https://github.com/difra100/Link_Prediction_with_PIGNN_IJCNN.
11footnotetext: Department of Computer Science, Control and Management Engineering, Sapienza University of Rome, Rome, Italy.22footnotetext: Institute of Information Science and Technologies ”Alessandro Faedo” - ISTI-CNR, Pisa, Italy.§§footnotetext: Corresponding author, difrancesco@diag.uniroma1.it.
I Introduction
Graph Neural Networks (GNNs) work as feature extractors that can be trained to perform some typical tasks on graphs, such as link prediction, and node or graph classification.
Among these, link prediction consists of computing the probability of a link between two nodes.
Most GNNs rely on the message-passing formalism [1]. In heterophilic graphs, where connected nodes tend to have different labels, Message-Passing Neural Networks (MPNNs) may struggle in the classification task as they tend to generate similar representations for adjacent nodes [2], an issue commonly known as over-smoothing [3]. There have been efforts to improve the performance of GNNs on heterophilic graphs [4], however, these works are limited to the context of node classification. An example of this is the gradient flow framework (GRAFF) [5], which deals with heterophily through physics-inspired biases. Although GRAFF shows competitive performance in both heterophilic and homophilic graphs, existing research has focused solely on node classification tasks, not investigating its potential in link prediction, which has significant applications in several domains. This task becomes particularly complex for heterophilic graphs, where we may have a link between two nodes with dissimilar characteristics. Unlike homophilic graphs, where connected nodes are similar, the reason behind the connection of two dissimilar entities can be latent. To the best of our knowledge, link prediction under heterophily has only been discussed and brought to the community’s attention by [6].
We propose GRAFF-LP (GRAFF for Link Prediction), a link prediction framework built upon GRAFF for node classification [5]. We tested our model on newly introduced heterophilic datasets [7] to overcome the limitations of standard benchmarks. The contributions of our work are the following.
  1. 1.
    We are the first to propose a Physics-Inspired GNN for link prediction.
  2. 2.
    We demonstrate that GRAFF-LP can achieve competitive performance w.r.t. the other examined methods, showing consistent performance across graphs from different contexts and increasing size.
  3. 3.
    We propose a novel Physics-Inspired readout function that leads to consistent performance improvements for GRAFF-LP as well as other baseline models. Additionally, the new readout gives GRAFF-LP more transparency in its behavior at inference time.
  4. 4.
    We set a new link prediction baseline on a recently created collection of heterophilic graphs [7], originally designed for node classification. This baseline serves as a foundation for future work in this area. Enhancing the current, yet not well-explored, literature on link prediction under heterophily.
  5. 5.
    Unexpectedly, we show that most of the time, classic models do not struggle in predicting the connections between nodes of different classes, conversely with what happens with heterophilic node classification.
Figure 1: General overview of the link prediction pipeline.
II Related Works
Graph Neural Networks for Link Prediction. Link prediction methods can be divided into non-neural-based approaches, such as heuristics [8], and neural-based methods, such as GNNs. Although heuristics work well in specific cases, GNNs offer a general framework by learning both graph structure and content features simultaneously. Neural-based methods include node-based models like Graph Auto-Encoders [9] and subgraph-based paradigm, led by SEAL [10]. However, while more expressive, subgraph-based methods are inefficient when scaling the graph size. Additionally, [11] showed that when datasets are carefully split, the performance gap between node-based and subgraph-based models is not as significant. Our work builds upon the node-based paradigm.
Physics-Inspired vs. Physics-Informed. Physics-Informed (PI) neural networks incorporate physical priors to improve performance and generalizability [12]. Physics-Inspired (PIrd) networks are a subset of PI methods, where physical constraints are embedded in the architecture itself, acting as inductive biases [13].
Physics-Inspired Graph Neural Networks. PIrd GNNs incorporate physics principles directly into the model’s structure. Examples include GNNs based on gradient flows [5], reaction-diffusion equations [14], based on nonlinear controlled and damped oscillators [13]. While these approaches have been applied to node classification, they remain unexplored for link prediction. Our work provides the first perspective on PIrd biases in this setting. The extended discussion on related works is available in Appendix Extended Related Works.
III Preliminaries
Notation. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph, with $\mathcal{V}$ the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ be the set of edges. $|\mathcal{V}|=N$ is the number of nodes. We denote by $\Gamma(i)$ the neighborhood of the node $i$. $\mathbf{D}$ is the diagonal matrix in $\mathbb{R}^{N\times N}$, such that $D_{ii}=|\Gamma(i)|$. $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the features of node $i$ and $y_{i}$ its label. The node representations can be ordered in a unique matrix, which we refer to as the instance matrix $\mathbf{X}\in\mathbb{R}^{N\times d_{0}}$. $\mathbf{A}\in\{0,1\}^{N\times N}$ is the adjacency matrix, with $A_{ij}=1$ if nodes $i$ and $j$ are connected, and $A_{ij}=0$ otherwise. We distinguish $\mathbf{H}^{T}$ from $\mathbf{H}^{\top}$, as the hidden representation of the nodes $\mathbf{H}$ at the time-step $T$, and the transpose operation $\cdot^{\top}$.
Homophily measures. The homophily assumption in graphs refers to the tendency of similar nodes to be connected. An unambiguous similarity measure is missing in the literature. Those that have been used the most with GNNs are edge homophily $\xi_{edge}$ [15] and node homophily $\xi_{node}$ [2]. The former was more considered within the node classification benchmarks [4, 5], and can be computed as
$\xi_{edge}=\frac{|(i,j)\in\mathcal{E}:y_{i}=y_{j}|}{|\mathcal{E}|}$ (1)
Both $\xi_{edge}$ and $\xi_{node}$ rely on the labels associated with the nodes. As an example, Equation (1) measures the fraction of edges that connect nodes from the same class. Generally, if we record a low homophily, we consider the graph as heterophilic. Traditional homophily measures are unsuitable for cross-datasets comparison because they are sensitive to class numbers and sample balance [16, 7]. To address this, the adjusted homophily metric $\xi_{adj}$ was introduced in [16]:
$\xi_{adj}=\frac{\xi_{edge}-\sum_{k\in\mathcal{S}}\mathcal{D}_{k}^{2}/(2|\mathcal{E}|)^{2}}{1-\sum_{k\in\mathcal{S}}\mathcal{D}_{k}^{2}/(2|\mathcal{E}|)^{2}},$ (2)
$\mathcal{Y}=\{1,....,C\}$ is the set of possible labels associated with each node, and $\mathcal{D}_{k}=\sum_{i:y_{i}=k}D_{ii}$. This measure is comparable across graphs and upper-bounded by 1, but it lacks a lower bound and does not consider node features.
Graph Neural Networks as gradient flow. Let us consider an $N$-dimensional dynamic system evolving as $\dot{\mathbf{H}}(t)=F(\mathbf{H}(t))$, with $\mathbf{H}(t)\in\mathbb{R}^{N\times d}$ If there exists a function $E:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}$, s.t. $F(\mathbf{H}(t))=-\nabla E(\mathbf{H}(t))$, the evolution equation $\dot{\mathbf{H}}(t)$ is the gradient flow of the energy $E$. Gradient flows are useful for studying the underlying dynamics of the system, provided the knowledge of $E$. In our case, $N$ represents the graph nodes whose representations $\mathbf{H}(t)$ evolve through a GNN over time, and $E(\mathbf{H}(t))$, is an energy functional associated with the node representations. Let $GNN:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be an intermediate layer of a generic GNN. By treating the GNN layers as continuous time $t$ and defining $GNN(\mathbf{H}(t))=-\nabla E(\mathbf{H}(t))$, the evolution of the features through the GNN is described as the gradient flow of $E(\mathbf{H}(t))$. $E$ can be selected as the Dirichlet Energy $E^{dir}$:
$E^{dir}(\mathbf{H}(t))$ $:=\sum_{(i,j)\in\mathcal{E}}\|(\nabla\mathbf{H}(t))_{ij}\|^{2}$ (3)
where $(\nabla\mathbf{H}(t))_{ij}=\frac{\mathbf{h}^{t}_{j}}{\sqrt{D_{jj}+1}}-\frac{\mathbf{h}^{t}_{i}}{\sqrt{D_{ii}+1}}$ and $\mathbf{h}^{t}_{i}$ denotes the feature of node $i$ at time $t$. Poor performance on heterophilic graphs and over-smoothing are often linked to the Dirichlet energy of features decaying to zero as layers increase [17, 18]. Equation (3) shows that decreasing $E^{dir}$ brings adjacent nodes closer in feature space. Conversely, from what was commonly thought, [5] proved that linear graph convolutions with symmetric weights shared among layers can induce edge-wise attraction (repulsion) through their positive (negative) eigenvalues. This control mechanism effectively influences whether the features are smoothed or sharpened, making the model successfully handle node classification within heterophilic graphs. This was possible by defining layers as a gradient flow of a parametrized Dirichlet energy:
$E_{\theta}^{dir}(\mathbf{H}(t))$ $=\sum_{i}\langle\mathbf{h}^{t}_{i},\mathbf{\Omega}\mathbf{h}^{t}_{i}\rangle-\sum_{i,j}a_{ij}\langle\mathbf{h}_{i},\mathbf{W}\mathbf{h}^{t}_{j}\rangle$ (4)
$=\sum_{i}\langle\mathbf{h}^{t}_{i},(\mathbf{\Omega}-\mathbf{W})\mathbf{h}^{t}_{i}\rangle+\frac{1}{2}\sum_{i,j}\|\Theta_{+}(\nabla\mathbf{H})_{ij}\|^{2}$
$-\frac{1}{2}\sum_{i,j}\|\Theta_{-}(\nabla\mathbf{H}(t))_{ij}\|^{2},$
$\Theta_{+}$ and $\Theta_{-}$ depend on the positive and negative eigenvalues of the weight matrices, respectively. By applying Euler discretization to the gradient flow and replacing $\mathbf{H}(t)$ by $\mathbf{H}^{t}$:
$\mathbf{H}^{t+\tau}=\mathbf{H}^{t}+\tau(-\mathbf{H}^{t}\mathbf{\Omega}+\boldsymbol{\mathbb{A}}\mathbf{H}^{t}\mathbf{W}-\mathbf{H}^{0}\tilde{\mathbf{W}})$ (5)
$\tau$ is the integration step, $T=\tau L$ is the total integration time, $\mathbf{\Omega}$, $\mathbf{W}$ and $\tilde{\mathbf{W}}$ are the trainable matrices, that are symmetric and shared across the layers, and $\boldsymbol{\mathbb{A}}=\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-\frac{1}{2}}$, since self-loops are included in $\mathbf{A}$ and $\mathbf{D}$ (i.e. $\tilde{\mathbf{A}}=\mathbf{I}+\mathbf{A}$). Equation (5) takes the form of a residual network [19], where the interpretation of GNNs as gradient flows is referred to in the literature as PIrd GNN. For details on these derivations, see [5]. In this work, we take advantage of these results and build upon this architecture for link prediction.
IV Proposed Framework: GRAFF-LP
GRAFF-LP operates in a transductive setting, where the graph retains all the nodes both in training and inference. Figure 1 presents a general overview of our approach. The whole scheme is designed as a node-based method for link prediction [8] and consists of three different phases.
Encoding phase. This is the transition from $\mathbf{X}$ to $\mathbf{H}^{0}=F_{\text{enc}}(\mathbf{X})$ In the experiments, we used one linear layer followed by dropout for all the models taken into consideration.
Message-Passing phase. This phase is formalized by Equation (5). $L$, $\tau$ and $d_{h}$, are hyperparameters of our architecture. The latter is the dimensionality of the intermediate representation that is fixed across the layers to maintain the dynamic system interpretation. Following the non-linear gradient flow approach proposed by [5], we interleave layers with non-linear functions, specifically the Rectified Linear Unit (ReLU). Although this means the network is no longer a discretized gradient flow, it still preserves the physical interpretation of the weight matrices. We adopt this strategy due to its improved performance over the linear counterpart in node classification experiments [5]. The message-passing of GRAFF-LP is:
$\mathbf{H}^{t+\tau}=\mathbf{H}^{t}+\tau\sigma(-\mathbf{H}^{t}\mathbf{\Omega}+\boldsymbol{\mathbb{A}}\mathbf{H}^{t}\mathbf{W}-\mathbf{H}^{0}\tilde{\mathbf{W}}),$ (6)
where $\sigma(\cdot)$ is the ReLU operation. To enforce the symmetry in the trainable parameters, we follow the diagonally-dominant approach used by [5].
Decoding phase. Let $\mathbf{z}_{i}=\mathbf{h}^{L}_{i}$ be the output of the message passing phase for node $i$. We describe the probability that nodes $i$ and $j$ share a link as $\hat{y}$:
$\hat{y}(\mathbf{z}_{i},\mathbf{z}_{j})=F_{dec}(\mathbf{z}_{i,j}),\;\;\text{ with }\mathbf{z}_{i,j}=f(\mathbf{z}_{i},\mathbf{z}_{j})$ (7)
$f(\cdot)$ is the readout function associated with the link, which aggregates the features coming from two nodes. $F_{dec}$ is an MLP of $L_{MLP}$ layers with width $d_{MLP}$ and nonlinear activations (see Figure 1). We use two types of readout in our experiments:
[…]
The Hadamard readout is commonly used with node-based link predictors [8]. In this paper, we are the first to propose an alternative function $f_{g}$, which we believe is more compatible with the GRAFF backbone. As seen in Equation (4), gradient flow operates by minimizing or maximizing the squared norm of edge gradients, and $f_{g}$ can be directly related to them through the following relation:
$||(\nabla\mathbf{H}^{T})_{i,j}||^{2}=\sum_{d}(f_{g}(\mathbf{z}_{i},\mathbf{z}_{j}))_{d}.$ (10)
More precisely, the gradient flow operates by minimizing or maximizing the squared norm of the gradients multiplied by the terms $\Theta_{+}$ and $\Theta_{-}$; we preferred to define the readout function without including those terms to speed up the calculation of $f_{g}$. We use the Hadamard product of the gradients to avoid any bias on the edge directionality since we deal with undirected graphs. However, for directed graphs, we could simply have $f_{g}(\mathbf{z}_{i},\mathbf{z}_{j})=(\nabla\mathbf{H}^{T})_{i,j}$. We do not consider other readouts such as concatenation of $z_{i}$ and $z_{j}$, since this would provide a bias on directionality as well, we would lose the physics-inspired bias, and we would double $d_{MLP}$, negatively affecting the space and time complexity.
All the node-based methods that we implement in our experiments follow the scheme in Figure 1, more details on each implementation can be found in the code.
IV-A Complexity Analysis of GRAFF-LP
We now consider the space complexity in terms of the number of parameters. GRAFF-LP, due to its weight-sharing mechanism, maintains constant complexity w.r.t. $L$ and exhibits quadratic complexity w.r.t. the number of hidden dimensions, as the elements in $\mathbf{\Omega}$ and $\mathbf{W}$ are of size $d_{h}^{2}$, where $d_{h}$ is the hidden dimension. Since these matrices are symmetric, there is redundancy in the parameters, reducing the total number to approximately $\frac{1}{2}d_{h}^{2}$.
To summarize, the number of parameters for GCN scales as $\mathcal{O}(Ld_{h}^{2})$. Since GCN shares a similar message-passing as Equation (5), but it does not use weight sharing, and we set $\mathbf{\Omega}\equiv 0$. In the case of GRAFF-LP, both $\mathbf{W}$ and $\mathbf{\Omega}$ use weight sharing, and the overall complexity scales as $\mathcal{O}(d_{h}^{2})$. In terms of time complexity, GRAFF-LP does not have any specific advantage over other models. In Table V, we show the runtime analysis for all the models, reporting also the number of parameters.
V Experiments
In this Section, we outline our experimental setup and results to address the following research questions:
RQ1:
How the $f_{g}$ readout contribute to the GNNs performance?
RQ2:
Can GRAFF-LP induce attraction and repulsion among existing and non-existing edges, resembling what is observed at node-level in node classification?
RQ3:
How much class heterophily impacts models’ performances?
TABLE I: Dataset Information
Datasets $N$ $|\mathcal{E}|$ $d$ $|C|$ $\xi_{edge}$ $\xi_{adj}$
Amazon Ratings 24492 186100 300 5 0.38 0.14
Roman Empire 22662 65854 300 18 0.05 -0.05
Minesweepers 10000 78804 7 2 0.68 0.01
Questions 48921 307080 301 2 0.84 0.02
Tolokers 11758 519000 10 2 0.59 0.09
V-A Experimental Set-up
V-A1 Datasets
Table I presents graph statistics after conversion to undirected graphs, a standard GNN procedure. The datasets selected belong to a recent collection [7], which was proposed to enrich the current dataset availability for the GNN experimental setting under heterophily. They have not yet been applied to link prediction. We used four datasets, Amazon Ratings, Roman Empire, Minesweeper, Questions, and Tolokers. Additional details and dataset descriptions can be found in Appendix Datasets Description. These datasets differ in context, size, and structural properties [7].
We did not consider datasets from the WebKB [20, 2] collection, since they lead to unstable and statistically insignificant results as already noted in [7], also in our experiments all the performances were not comparable, making it impossible to understand what model was more effective w.r.t. the others.
V-A2 Baselines
We compare against several baselines, including a Multi-layer Perceptron (MLP) using only node features, and both node-based and subgraph-based methods.
GCN and GraphSAGE: Convolutional MPNNs [21]; for GraphSAGE, we evaluate both mean and max variants.
GAT: An attentional MPNN [21] that uses an attention mechanism for neighbor aggregation, particularly effective for heterophilic graphs.
ELPH: [22] A subgraph-based method for link prediction that avoids explicit subgraph computation, using a GCN-based feature extractor, though compatible with other MPNNs.
NCNC: [23] A link prediction method, based on structural features as ELPH, but these are related to higher-order common neighbor information. This method is currently state-of-the-art in a recent link prediction benchmark [11].
We also intended to test Disenlink [6], a model designed for link prediction under heterophily, but the official implementation prevented its use on our datasets, because of their size. This was due to an inefficient implementation and use of the adjacency matrix.
Further experimental details are available in Appendix Experimental Set-up: Additional Details.
V-B Results
V-B1 Effectiveness
To evaluate model performance, we use the Area Under the Receiver Operating Characteristic (AUROC). Table II summarizes the results across all datasets for both readout functions, $f_{h}$ and $f_{g}$. The results with $f_{h}$ are related to the best configuration of hyperparameters with $f_{h}$. Concerning $f_{g}$, we just took the same configuration found with $f_{h}$ and trained the models again via $f_{g}$. Performance metrics are averaged over 10 random seeds. We adopted a single split since the graphs are sufficiently large and less sensitive to high variance in the data. The statistical significance of the results is computed through the Wilcoxon test [24], via the accuracy of positive and negative edges in the test set; details about the procedure can be found in the code.
With $f_{h}$, GRAFF-LP leads in 3 out of 4 datasets and in 2 out of 4 using $f_{g}$, ranking consistently in the top 2 across all datasets, indicating adaptability not seen with the other models. These aspects also apply to Tolokers, whose results and comments are reported in Appendix Additional Results.
On the Roman Empire dataset, which resembles a chain-like graph, GRAFF-LP achieves the highest AUROC in both setups, with a neat advantage w.r.t. other baselines. ELPH, and NCNC also present a wide gap w.r.t node-based methods, but they have access to the subgraph features that are a clear advantage to perform link prediction in such a structure. This advantage and superiority of ELPH and NCNC are found with Minesweeper, where the graph is a regular grid, and link prediction can be easily solved by understanding the grid structure.
In both cases, the physics-inspired bias seems to let the model use the structure properly and achieve state-of-the-art performance. According to our hyperparameter optimization, ELPH fails to obtain competitive performance in Amazon Ratings and Questions. We relate this gap to the limited depth of ELPH (which is up to 3 layers). NCNC instead can reach competitive performance in Amazon Ratings, as well as Questions. Despite this, it never surpasses GRAFF-LP. We conclude that subgraph-based methods do not perform better than GRAFF-LP on class heterophilic datasets. This implies that structural features are not critical or determining to perform well in these settings.
Another observation is that GCN, both in Amazon Ratings and Questions, is among the top-scoring models, and in particular gets the closest to GRAFF-LP. This is not unexpected, since they share a similar message-passing and expressivity [25, 5]. They mainly differ in the weight design. We associate these similarities to the same phenomenon observed by [5], who observed that in homophilic node classification, the two models perform the same. We conjecture in this case that for Amazon Ratings and Questions the edge-wise repulsion is not required, thus GCN can afford similar results. We also report the results of the MLP. The MLP results consistently lag behind the GNN models, emphasizing the need for graph representation learning for effective link prediction in these new datasets. For example, in Minesweeper, MLP struggles to infer missing links due to its lack of graph context and the dataset’s grid structure.
[…]
Datasets Amazon Ratings Roman Empire Minesweeper Questions
Models $f_{h}$ $f_{g}$ $f_{h}$ $f_{g}$ $f_{h}$ $f_{g}$ $f_{h}$ $f_{g}$
MLP 66.80 $\pm$ 9.5* 70.18 $\pm$ 0.1* 64.74 $\pm$ 1.7* 64.60 $\pm$ 1.1* 59.19 $\pm$ 1.8* 47.26 $\pm$ 3.8* 77.44 $\pm$ 0.7* 73.92 $\pm$ 3.9*
GCN 93.97 $\pm$ 0.8* 98.90 $\pm$ 0.2* 51.32 $\pm$ 1.6* 56.28 $\pm$ 4.1* 94.44 $\pm$ 1.0* 98.56 $\pm$ 0.3* 97.63 $\pm$ 0.1*** 97.56 $\pm$ 0.1
SAGE 65.58 $\pm$ 22.5* 69.24 $\pm$ 5.5* 65.17 $\pm$ 1.4* 66.57 $\pm$ 3.6* 97.99 $\pm$ 0.7* 96.40 $\pm$ 2.1* 91.48 $\pm$ 1.4* 93.20 $\pm$ 0.9*
GAT 60.31 $\pm$ 6.0* 72.22 $\pm$ 6.1* 71.33 $\pm$ 1.4* 73.85 $\pm$ 3.4* 95.05 $\pm$ 2.8* 98.23 $\pm$ 0.7* 78.66 $\pm$ 9.7* 69.81 $\pm$ 3.6*
ELPH 55.89 $\pm$ 12.5* 55.89 $\pm$ 12.5* 87.01 $\pm$ 1.1* 87.01 $\pm$ 1.1* 99.88 $\pm$ 0.2 99.88 $\pm$ 0.2 82.13 $\pm$ 10.2* 82.13 $\pm$ 10.2*
NCNC 98.20 $\pm$ 1.3 98.20 $\pm$ 1.3* 86.73 $\pm$ 6.5* 86.73 $\pm$ 6.5* 95.42 $\pm$ 11.4* 95.42 $\pm$ 11.4* 94.61 $\pm$ 0.5* 94.61 $\pm$ 0.5*
GRAFF-LP 98.69 $\pm$ 0.4 99.47 $\pm$ 0.1 98.23 $\pm$ 0.8 99.34 $\pm$ 0.4 99.01 $\pm$ 0.4 99.41 $\pm$ 0.2 97.64 $\pm$ 0.0 97.54 $\pm$ 0.1
In Appendix -D, we provide examples of GRAFF-LP performing on link prediction under homophily, to show that our approach can reach competitive as well as state-of-the-art performance, also under homophily.
In Table III, we show the relative percentage improvement we obtained using $f_{g}$ as the readout method. We did not include ELPH and NCNC, since their official derivation is based on Hadamard or other readouts.
TABLE III: Percentage Increase of Models Across Datasets, when going from $f_{h}$ to $f_{g}$.
Model A. Ratings R. Empire Minesweeper Questions
MLP +4.48% 0% -20.34% -3.90%
GCN +5.32% +9.80% +5.32% 0%
SAGE +4.55% +3.08% -2.04% +2.20%
GAT +20% +4.23% +3.16% -11.39%
GRAFF-LP +0.70% +1.02% 0% 0%
V-B2 Physics-Inspired Link Prediction
[5] proposed the GRAFF architecture and showed the attraction and repulsion behavior among adjacent nodes with different labels. This is expressed in Equation (4). In the link prediction scenario, we believe that the same behavior can be induced among two nodes that, even though they seem different, should attract themselves, and vice versa, where two nodes are similar but do not present any interaction. Since we build upon the same framework of [5], we expect the edge gradients associated with $\Theta_{+}$ to become smaller, because of attraction in the feature space, and the edge gradients associated with $\Theta_{-}$ to become larger in the feature space because of repulsion. In this way, we can deduce that GRAFF-LP can induce attraction and repulsion in the same fashion as in GRAFF for node classification, and we would answer positively to RQ2. To understand whether this phenomenon happens, we introduce a novel metric, namely the gradient separability ($GS$). Let us assume we have a set of edge gradients =$\nabla=\{(\nabla\mathbf{H}^{t})_{i,j}|0\leq t\leq T\wedge(i,j)\in\mathcal{E}_{t}\}$, where $\mathcal{E}_{t}=\{\mathcal{E}_{pos}\cup\mathcal{E}_{neg}\}$, is the set of evaluation edges, for example the test edges, which comprises both positive edges $\mathcal{E}_{pos}$, and negative edges $\mathcal{E}_{neg}$. Roughly speaking, $\nabla$ contains the edge gradients of positive and negative test edges computed at each layer of the model. This is done only for message-passing layers, excluding those in the Encoding and Decoding phases, to isolate the effect of the message-passing. We specify the squared norm of the positive (or negative) edge gradients at time t as follows:
$\nabla^{t}_{pos}$ =$=\{||(\nabla\mathbf{H}^{t})_{i,j}||^{2}|(i,j)\in\mathcal{E}_{pos}\}$ (11)
$\nabla^{t}_{neg}$ =$=\{||(\nabla\mathbf{H}^{t})_{i,j}||^{2}|(i,j)\in\mathcal{E}_{neg}\}$ (12)
More formally, by minimizing (4), we expect the squared norm of the positive (negative) gradients to decrease (increase). To measure this, we take advantage of the AUROC metric, considering the positives as class 0, while the negatives as class 1. Through the AUROC, we evaluate how much these two classes are separated based on their scores, regardless of any threshold. In particular if we consider as ground truth a stack of 1’s and 0’s for negatives and positives $\mathcal{S}_{\mathcal{E}_{t}}=\{\mathbb{1}_{\{(i,j)\in\mathcal{E}_{neg}\}}|\forall(i,j)\in\mathcal{E}_{t}\}$, $GS$ at time $t$ is computed as
$GS^{t}=AUC(\mathcal{S}_{\mathcal{E}_{t}},\{\nabla^{t}_{pos}\cup\nabla^{t}_{neg}\}).$ (13)
This way, we can monitor $GS^{t}$ over time to understand the trend of the edge gradients. We can also distinguish the edges connecting nodes of the same class, and those connecting edges from different ones.
Let us define $\nabla^{t}_{pos}=\{\nabla^{t}_{pos,hm}\cup\nabla^{t}_{pos,ht}\}$, $\nabla^{t}_{neg}=\{\nabla^{t}_{neg,hm}\cup\nabla^{t}_{neg,ht}\}$, as the composition of the edge gradients coming from homophilic edges $\mathcal{E}_{hm}=\{(i,j)\in\mathcal{E}\wedge y_{i}=y_{j}\}$, and heterophilic ones $\mathcal{E}_{ht}=\{(i,j)\in\mathcal{E}\wedge y_{i}\neq y_{j}\}$, according to this distinction we can evaluate the subset of the edges based on their class labels. The gradient separability can be written accordingly:
$GS^{t}_{\mathcal{U},\mathcal{V}}=AUC(\mathcal{S}_{\mathcal{U}\cup\mathcal{V}},\{\nabla^{t}_{pos,\mathcal{U}}\cup\nabla^{t}_{neg,\mathcal{V}}\})$ (14)
$GS^{t}_{ht,hm}$ measures the ability to classify as 0 the positive heterophilic edges, and as 1 the negative homophilic edges. In other words, it measures how much we can separate the positive and the negative edges based on the squared norm of their edge gradient. We provide a visual example in Figure 2, where we consider 1 on the y-axis as the homophilic edges, while the 0 y-axis refers to the heterophilic edges. We visualize a total of 50 samples, but the scores refer to those computed within the whole test set. These are produced from a fully-trained GRAFF-LP, monitoring $GS^{t}_{hm,hm}$ and $GS^{t}_{ht,ht}$. Figure 3 also includes the distributions of the squared norm gradients to better understand how they separate.
(a) $GS^{0}_{hm,hm}$, $GS^{0}_{ht,ht}$
(b) $GS^{3}_{hm,hm}$, $GS^{3}_{ht,ht}$
(c) $GS^{6}_{hm,hm}$, $GS^{6}_{ht,ht}$
(d) $GS^{9}_{hm,hm}$, $GS^{9}_{ht,ht}$
Figure 2: $||(\nabla\mathbf{H}^{t})_{i,j}||^{2}$ evolution with a fully-trained 9-layer GRAFF-LP via $f_{g}$ on Minesweeper.
Figure 3: Minesweeper: Edge Gradients Distribution after each message-passing phase, for a 9-layer GRAFF-LP.
These visualizations, along with those in Appendix -E, provide qualitative and quantitative evidence that GRAFF-LP induces edge attraction and repulsion via $f_{g}$ as the nodes go through the network’s layers. We answer positively to RQ2.
In Table IV, instead we show $GS^{T}$, which is the total gradient separability computed after the last message-passing layer, for each model and averaged across 10 random seeds. We report the results and comments about Tolokers in Appendix -C. In this analysis, we do not use any statistical test, since we do not want to assess which model has the highest $GS^{T}$; we want to understand what models have a sufficiently high value of $GS^{T}$ to acknowledge their ability to separate edge gradients. We consider $GS^{T}>90\%$ as evidence for the model to be able to distinguish the nature of the edges, based uniquely on the squared norm of the gradients. The most interesting aspect of $GS^{T}$ is the consistent increase that we record for GRAFF-LP, which we associate with the capability of this model to understand the negative and positive edges based on the evolution of their gradients. Indeed, obtaining $GS^{T}=100\%$ implies that the edges can be classified as positive or not, based solely on $\{\nabla^{T}_{pos},\nabla^{T}_{neg}\}$, which are available even before the Decoding phase. According to our results, GRAFF-LP allows us to do so even though it is not trained explicitly to do it. Equation (4) lets the model minimize the squared norm of the gradients, but it refers to the message-passing edges, not to the unseen edges that we want to predict. In the Roman Empire experiments, when we go from $f_{h}$ to $f_{g}$ the AUROC improvement is limited to $+1.02\%$ (see Table III), but when it comes to gauging the gradient separability difference, we have a $+65.52\%$. Since the only difference in the two models is the readout, we affirm that $f_{g}$ induces the model to learn to separate the edges following a strategy driven by the gradients. The only dataset where GRAFF-LP seems not to catch this behavior is Questions, where $GS^{T}$ is close to 0. Anyway, this is not an issue, since the AUROC is faulty when it approaches 50%, in our case, it means that the negative edges decrease, rather than the positive ones. For this reason, we conjecture that GRAFF-LP is learning the opposite behavior. The same happens to GCN, which is even closer to 0 than GRAFF-LP since it reaches down to $GS^{T}=14.32\%$. Surprisingly, not only is our model able to learn to separate the edge gradients, but also the other node-based methods, except for the MLP, ELPH, NCNC. Nonetheless, NCNC scores $GS^{T}>90$ in Minesweeper. Despite this, we recognize GRAFF-LP as the approach that consistently shows this behavior; we can conclude that this is due to the PIrd message-passing and readout schemes. ELPH reaches state-of-the-art performance in Minesweeper, leaving GRAFF-LP as the second best performing model, but our model offers more transparency in the model’s behavior, as confirmed by $GS^{T}$. These results describe the full contribution of our newly introduced readout function $f_{g}$ answering RQ1. Specifically, $f_{g}$ can enhance the prediction performance (see Table III), and also the model’s transparency at the inference phase, as illustrated in this Section.
TABLE IV: Comparison of $GS^{T}$, when the model is trained with $f_{h}$ or $f_{g}$. The score that is closest to 100 or 0 is highlighted in bold. We report the percentage variation between the results of $f_{h}$ and $f_{g}$ as $\Delta$.
Model Readout Type Amazon Ratings Roman Empire Minesweeper Questions
$GS^{T}$ $\Delta$ $GS^{T}$ $\Delta$ $GS^{T}$ $\Delta$ $GS^{T}$ $\Delta$
MLP $f_{h}$ 48.49 ± 14 - 35.97 ± 0.42 - 63.56 ± 1.6 - 47.65 ± 9.93 -
$f_{g}$ 30.04 ± 0.06 -37.5% 36.92 ± 1.7 +2.78% 62.04 ± 2.4 -3.13% 43.18 ± 4.6 -10.42%
GCN $f_{h}$ 71.9 ± 2 - 35.78 ± 0.2 - 76.7 ± 2.1 - 15.57 ± 2.5 -
$f_{g}$ 86.32 ± 3.9 +19.44% 35.79 ± 0.2 0% 94.44 ± 1.1 +22.08% 14.32 ± 1.9 -12.5%
SAGE $f_{h}$ 30.12 ± 0.05 - 37.26 ± 1.5 - 77.69 ± 1.6 - 63.31 ± 1.7 -
$f_{g}$ 30.11 ± 0.06 0% 40.56 ± 3.6 +10.81% 68.81 ± 1.4 -11.54% 63.41 ± 0.9 0%
GAT $f_{h}$ 33.45 ± 4.3 - 42.66 ± 4.8 - 93.22 ± 1.7 - 39.14 ± 3.9 -
$f_{g}$ 31.13 ± 2.5 -6.06% 46.24 ± 2.9 +6.98% 97.45 ± 1.01 +4.3% 37.42 ± 1.07 -5.13%
ELPH $f_{h}$ 46.76 ± 7.3 - 36.16 ± 0.11 - 73.44 ± 5.7 - 31.92 ± 8.08 -
NCNC $f_{h}$ 30.49 ± 0.04 - 36.55 ± 0.23 - 91.31 ± 0.99 - 37.44 ± 1.08 -
GRAFF-LP $f_{h}$ 94.52 ± 2.04 - 58.46 ± 2.7 - 93.59 ± 2.13 - 26.74 ± 3.71 -
$f_{g}$ 96.43 ± 0.9 +1.05% 95.96 ± 2.4 +65.52% 98.10 ± 0.32 +4.26% 18.20 ± 3.5 -33.33%
V-B3 Runtime Analysis
We briefly show some results about the time and space complexities that these baselines have. We already discussed space complexity in terms of parameters, but in this Section, we report the real-world values associated with the experiments. In Table V, we show the number of parameters of each model and the inference time, computed on the test edges for 10 epochs. All the models have $L=3$, $L_{MLP}=1$, and $d_{h}=d_{MLP}=64$, to make a fair comparison across models. What we notice is that according to our discussion on complexity, GRAFF-LP is the lightest model, and also comparable with the other node-based methods such as GCN, GAT and GraphSAGE. ELPH, which is a subgraph-based method has the highest inference time, as expected. While NCNC, even though it is known to be more efficient than ELPH, and more similar to the node-based methods, it still presents a significant gap in terms of inference, due to the computation of structural features. Furthermore, it leads to out-of-memory issues in Tolokers. In Appendix Extended Runtime Analysis, we report additional results on runtime analysis.
TABLE V: Comparison of model performance across datasets, showing the number of parameters and runtime (in seconds) for each model. The inference time is averaged across 10 trials.
Model Amazon Ratings Minesweeper
Parameters Runtime (s) Parameters Runtime (s)
MLP 32256 $0.0909\pm 0.01$ 13504 $0.0512\pm 0.01$
GCN 31680 $0.118\pm 0.01$ 12928 $0.0832\pm 0.01$
GAT 32064 $0.1079\pm 0.01$ 13312 $0.0727\pm 0.01$
SAGE 43968 $0.0919\pm 0.02$ 25216 $0.0818\pm 0.01$
ELPH 40542 $0.8751\pm 0.05$ 21790 $0.4563\pm 0.03$
NCNC 27584 $0.1935\pm 0.02$ 8832 $0.1143\pm 0.01$
GRAFF-LP ($f_{h}$) 23617 $0.1072\pm 0.01$ 4865 $0.0756\pm 0.01$
GRAFF-LP ($f_{g}$) 23617 $0.1021\pm 0.01$ 4865 $0.0796\pm 0.01$
V-B4 Does Class Heterophily Impact GNN performance?
In this paper’s premises, we questioned the absence of explicit methods that tackle heterophily in link prediction. This is a gap in the literature that is worth studying since we have seen in the previous Sections, that in some datasets, even subgraph-based approaches fail to achieve competitive performance. Other than the mere performance, in the literature is pointed out the complicated nature of the task [6], namely, how can the latent factors that connect two entities be understood? In this paper, we focused on datasets that are known to be heterophilic for node classification. Are they representative enough for link prediction under heterophily, even if $\xi_{adj}$ is determined only from class labels? Do simple GNN baselines struggle to learn from these graphs, in the same fashion as node classification? In other words, can they learn to predict existing(non-existing) links when two entities have different(same) classes? These concerns are related to RQ3. To answer it, we measure the AUROC that we obtain when we try to classify all the possible class mixes, namely when the positive or negative edges are homophilic or heterophilic. In this way, it is possible to measure 4 different AUROCs that are representative of the model’s ability to separate edges of different natures.
Let us define $AUC_{\mathcal{U},\mathcal{V}}$ as the ability of a binary classifier to correctly distinguish the sets $\mathcal{U}$ and $\mathcal{V}$ as existing or non-existing edges. Here, the ground truth is that $\mathcal{U}$ contains positives and $\mathcal{V}$ the negatives. The nature of these can be homophilic or heterophilic, namely $\mathcal{U},\mathcal{V}\in\{hm,ht\}$. In Figure 4, we take into consideration Roman Empire as an example of a dataset where the models fail to achieve high performance, and we show that is not due to heterophilic edges (i.e. $AUC_{ht,hm},AUC_{ht,ht}$), since we have poor performance also on the homophilic ones $AUC_{hm,ht},AUC_{hm,hm}$. In Appendix -F, we provide additional evidence of this behavior, which also manifests when models reach high performance. Generally, we understand that the performance does not vary significantly, and the models’ rankings remain consistent. We conclude that models that achieve competitive performances in homophilic link prediction do not perform as well in our benchmark, but this is not due to class heterophily. This result answers negatively to RQ3 and highlights the need for new homophilic measures that can better identify datasets representative of link prediction under heterophily, because of its relevance in several applications.
Figure 4: Roman Empire: Ability of the model to predict homophilic edges or heterophilic both as negatives or positives.
VI Conclusions
This paper presented GRAFF-LP, a link prediction model built upon GNNs and driven by a PIrd bias. Our method is designed to handle heterophilic graphs, enriching the current literature available on the topic, which is often confined to node classification. We show that GRAFF-LP outperforms the other baselines in 3 out of 4 datasets. We also propose $f_{g}$, a PIrd readout function based on the edge gradients, that, without any tuning, is able to improve the GRAFF-LP performance. Surprisingly, we observe that other models also benefit from $f_{g}$, but GRAFF-LP consistently ranks among the top 2. To better understand the effect of $f_{g}$, we study the evolution of the squared norm of the edge gradients via a novel metric named $GS$, and we surprisingly found that, under the lens of this norm, the GRAFF-LP’s message-passing separates the positive and negative edges. GRAFF-LP showcases this behavior more consistently than the other baselines, favored by the attraction and repulsion bias of the gradient flow interpretation. We also brought some evidence that the baselines do not struggle to perform link prediction because of class heterophily. This result exposes the need for new measures that help identify challenging scenarios to advance our knowledge of link prediction under heterophily. Future works should address the aspects related to new heterophily metrics based on the input features since they are not task-dependent. Moreover, we plan to further understand how GRAFF-LP can learn the edge gradient separation, even though it was not trained explicitly to do so. Then, it would be interesting to adapt other PIrd GNNs or develop new ones specifically suited for link prediction.